# Optimal Reward Labeling: Bridging Offline Preference and Reward-Based Reinforcement Learning

### v1

@@ 6 Conclusion and Limitation
Therefore, the prediction loss is monotonically increasing on $r^{i}_{\sigma^{i},t}$ and monotonically decreasing on $r^{i}_{\bar{\sigma}^{i},t}$ for all $i\in[N]$ and $t\in[T]$. Given that the rewards are bound in $[-1,1]$, the optimal reward that achieves minimal prediction loss is $r^{i}_{\sigma^{i},t}=1$ and $r^{i}_{\bar{\sigma}^{i},t}=-1$ for all $i\in[N]$ and $t\in[T]$.
Appendix B Analyze for Combining ORL with Model-Free Standard Offline RL Algortihms
In a model-free offline RL algorithm, training a pair of actor and critic models is common [Fujimoto and Gu, 2021, Cheng et al., 2022]. The actor model is usually a policy that should be optimal, and the critic model is usually a Q function that represents the long-term rewards for taking an action at a state. In the reward-based setting, the critic model utilizes the reward signals for update through a quantity called ‘Bellman Loss’. Given a tuple $(s,a,r,s^{\prime})$, a critic $Q$, and an actor $\pi$, the Bellman loss of the critic on the data tuple is defined as $|Q(s,a)-(r+\gamma\cdot\max_{a^{\prime}}Q(s^{\prime},a^{\prime})|$. Bellman loss acts as a criterion for evaluating the quality of the $Q$ function.
Here, we analyze the case when applying a model-free algorithm to our framework. In a model-free offline RL algorithm, training a pair of actor and critic models is common [Fujimoto and Gu, 2021, Cheng et al., 2022]. The actor model is usually a policy that should be optimal, and the critic model is usually a Q function that represents the long-term rewards for taking an action at a state. In the reward-based setting, the critic model utilizes the reward signals for update through a quantity called ‘Bellman Loss’. Given a tuple $(s,a,r,s^{\prime})$, a critic $Q$, and an actor $\pi$, the Bellman loss of the critic on the data tuple is defined as $|Q(s,a)-(r+\gamma\cdot\max_{a^{\prime}}Q(s^{\prime},a^{\prime})|$. Bellman loss acts as a criterion for evaluating the quality of the $Q$ function.
In a model-free algorithm, the reward signals influence the update of the $Q$ function through the gradient of the Bellman loss. In the case of optimal reward labels, the Bellman loss, as well as its gradient, are given in Eq 3.
$L_{3}=$ $\sum_{(s,a,s^{\prime})\in\text{Chosen}}|\hat{Q}(s,a)-(1+\gamma\cdot\max_{a^{\prime}}\hat{Q}(s^{\prime},a^{\prime}))|+$ (3)
$\sum_{(s,a,s^{\prime})\in\text{Rejected}}|\hat{Q}(s,a)-(-1+\gamma\cdot\max_{a^{\prime}}\hat{Q}(s^{\prime},a^{\prime}))|$
$\frac{\partial L_{3}}{\partial w}=$ $\sum_{(s,a,s^{\prime})\in\text{Chosen}}\frac{\partial\hat{Q}(s,a)-\gamma\cdot\max_{a^{\prime}}\hat{Q}(s^{\prime},a^{\prime}))}{\partial w}-$
$\sum_{(s,a,s^{\prime})\in\text{Rejected}}\frac{\partial\hat{Q}(s,a)-\gamma\cdot\max_{a^{\prime}}\hat{Q}(s^{\prime},a^{\prime}))}{\partial w}$
In the derivation of the gradient, we require $\hat{Q}(s,a)-\gamma\cdot\max_{a^{\prime}}\hat{Q}(s^{\prime},a))$ to be bounded as it represents the reward corresponding to the $Q$ function.
If one wants to modify a model-free reward-based algorithm to work with preference signals directly, a straightforward way is to replace the Bellman error with the prediction error of the $Q$ function on the preference signals. Based on this method, Hejna and Sadigh [2024] developed the IPL algorithm for PBRL by modifying the IQL algorithm [Kostrikov et al., 2021], a famous method for the standard RL problem. Specifically, the underlying reward function for a $Q$ function is given by $\hat{r}(s,a)=\hat{Q}(s,a)-\gamma\cdot\max\hat{R}(s^{\prime},a^{\prime})$. This underlying reward function can be used to predict the probability that the chosen trajectory is preferred to represent the prediction of the $Q$ function. The prediction loss of the $Q$ function on the preference labels and its gradient are given in Eq 4.
$L_{4}=$ $F(\sum_{(s,a,s^{\prime})\in\text{Chosen}}(\hat{Q}(s,a)-\gamma\cdot\max_{a^{\prime}}\hat{Q}(s^{\prime},a)))-$ (4)
$\sum_{(s,a,s^{\prime})\in\text{Rejected}}(\hat{Q}(s,a)-\gamma\cdot\max_{a^{\prime}}\hat{Q}(s^{\prime},a)))$
$\frac{\partial L_{4}}{\partial w}=$ $F^{\prime}(\sum_{(s,a,s^{\prime})\in\text{Chosen}}(\hat{Q}(s,a)-\gamma\cdot\max_{a^{\prime}}\hat{Q}(s^{\prime},a)))-$
$\sum_{(s,a,s^{\prime})\in\text{Rejected}}(\hat{Q}(s,a)-\gamma\cdot\max_{a^{\prime}}\hat{Q}(s^{\prime},a)))$
$\cdot$ $(\sum_{(s,a,s^{\prime})\in\text{Rejected}}\frac{\partial\hat{Q}(s,a)-\gamma\cdot\max_{a^{\prime}}\hat{Q}(s^{\prime},a))}{\partial w}-$
$\sum_{(s,a,s^{\prime})\in\text{Chosen}}\frac{\partial\hat{Q}(s,a)-\gamma\cdot\max_{a^{\prime}}\hat{Q}(s^{\prime},a))}{\partial w})$
Again, note that $F^{\prime}$ is always negative, as the link-loss function $F$ is monotonically decreasing. Comparing the two equations, we find that the directions of the gradients in both cases are the same. If the link-loss function $F$ is linear, then the gradients in both cases are only different by a constant multiplier. In other words, under this condition, both approaches utilize the preference labels in the same way for updating the critic model.
Appendix C Reward Labeling on Dataset of Overlapped Trajectories
Here, we consider the case where the same state action appears in different trajectories, and the reward labels are generated from the same reward model. As we discussed in Section 4, in this case, the labeling technique from Lemma 4.1 is no longer optimal in this case. In experiments, we apply the ORL framework on the datasets whose trajectories do not overlap with each other and the datasets of the opposite case, which are the ones in the main paper. The results in Fig 6 show that our ORL framework has equally high performance in both cases.
Figure 6: Comparison between the cases where the trajectories overlap with each other or not in the dataset. Our ORL framework achieve similar high learning performance in both cases.
Experimental Statistical Significance.
In Table 1, 2, 3, we show the statistical analysis results for the ablation studies our experiments. For clarity, we use a representative number to denote the learning efficiency of a method based on its training log. The number is the average of the performance of the learned policies from the last $10$ checkpoints. The results here show that the variance is relatively small in our experiments.
Halfcheetah Medium Expert Hopper Medium Replay
100% 50% 20% 10% 100% 50% 20% 10%
Oracle 85.0 $\pm$ 6.5 87.5 $\pm$ 2.8 89.3 $\pm$ 3.8 83.7 $\pm$ 7.1 76.7 $\pm$ 8.0 55.8 $\pm$ 37.4 13.9 $\pm$ 7.0 4.8 $\pm$ 1.8
ORL 91.9 $\pm$ 2.0 89.7 $\pm$ 2.8 76.9 $\pm$ 6.0 64.8 $\pm$ 5.7 32.0 $\pm$ 13.7 29.2 $\pm$ 6.5 5.6 $\pm$ 4.2 7.4 $\pm$ 5.9
RM 43.1 $\pm$ 0.27 43.4 $\pm$ 0.65 43.9 $\pm$ 0.81 48.8 $\pm$ 2.8 3.5 $\pm$ 3.4 4.0 $\pm$ 3.8 7.9 $\pm$ 3.4 4.4 $\pm$ 4.8
IPL 35.4 $\pm$ 2.7 32.9 $\pm$6.1 34.4 $\pm$ 3.9 32.4 $\pm$ 4.5 8.63 $\pm$1.8 13.8 $\pm$ 7.8 10.2 $\pm$ 5.0 8.3 $\pm$ 4.9
Table 1: Performance comparison of the same learning method on datasets of different sizes. The number is the average of the performance of the learned policies from the last $10$ checkpoints.
Halfcheetah-Medium-Expert IQL CQL COMBO
Oracle 91.3 $\pm$ 3.4 86.0 $\pm$ 5.9 97.8 $\pm$ 0.91
ORL 85.4 $\pm$ 4.4 91.9 $\pm$ 2.0 100.0 $\pm$ 3.3
Table 2: Performance comparison of the different learning method on the same dataset.
Halfcheetah-Medium-Expert Hopper-Medium-Replay
Oracle 86.0 $\pm$ 5.9 76.7 $\pm$ 8.0
ORL-Single 91.9 $\pm$ 2.0 32.0 $\pm$ 13.7
ORL-Multi 90.8 $\pm$ 4.4 31.1 $\pm$ 5.7
RM-Single 43.1 $\pm$ 0.27 3.5 $\pm$ 3.4
RM-Multi 43.5 $\pm$ 0.64 4.7 $\pm$ 3.0
Table 3: Performance comparison of the same learning method on datasets of different types of preference labels.
Compute resource. Our experiments are basic RL experiments for continuous control problems that require limited computing resources. Specifically, we run our experiments on a machine with 1. one CPU (11th Gen Intel(R) Core(TM) i9-11900K @ 3.50GHz) and 2. one GPU (NVIDIA RTX A4000 (16114MiB VRAM))

### v3

@@ Appendix B Proof
Therefore, the prediction loss is monotonically decreasing on $r^{i}_{\sigma^{i},t}$ and monotonically increasing on $r^{i}_{\bar{\sigma}^{i},t}$ for all $i\in[N]$ and $t\in[T]$. Given that the rewards are bound in $[-1,1]$, the optimal reward that achieves minimal prediction loss is $r^{i}_{\sigma^{i},t}=1$ and $r^{i}_{\bar{\sigma}^{i},t}=-1$ for all $i\in[N]$ and $t\in[T]$.
B.2 Proof for Theorem 4.5 and A.3
In the first case, when there is no overlap between trajectories, by Lemma 4.2, the optimal reward labels are the binary reward labels in Eq 2. For the two reward modeling processes, the optimal solutions are the reward models that output exactly the same binary reward labels. For the two Q-learning processes, the optimal solutions are the Q functions whose corresponding reward models output exactly the same binary reward labels.
In the second case, recall that the rewards functions and the corresponding reward functions of the Q functions are bounded. We can rewrite $\mathcal{L}_{1}$ and $\mathcal{L}_{3}$ as
$\mathcal{L}_{1}$ $=2\cdot T-\sum_{(s,a)\in\tau^{i}_{1}}\widehat{\mathcal{R}}(s,a)|+\sum_{(s,a)\in\tau^{i}_{2}}|\widehat{\mathcal{R}}(s,a)+1|$ (3)
$\mathcal{L}_{3}$ $=2\cdot T-\sum_{(s,a,s^{\prime})\in\tau^{i}_{1}}(\hat{Q}(s,a)-\gamma\cdot\max_{a^{\prime}}\hat{Q}(s^{\prime},a^{\prime}))+\sum_{(s,a,s^{\prime})\in\tau^{i}_{2}}(\hat{Q}(s,a)-\gamma\cdot\max_{a^{\prime}}\hat{Q}(s^{\prime},a^{\prime}))$
By comparing that with $\mathcal{L}_{2}$ and $\mathcal{L}_{4}$, we get the statements in the second case for both theorems.
In the third case, we check the gradients of the optimization goals at each trajectory pair:
$L_{1}$ $=\sum_{(s,a)\in\tau^{i}_{1}}|1-\hat{R}(s,a)|+\sum_{(s,a)\in\tau^{i}_{2}}|\hat{R}(s,a)+1|$ (4)
$=\sum_{(s,a)\in\tau^{i}_{1}}-\hat{R}(s,a)+\sum_{(s,a)\in\tau^{i}_{2}}\hat{R}(s,a)+2\cdot T$
$\frac{\partial L_{1}}{\partial w}$ $=\sum_{(s,a)\in\tau^{i}_{1}}-\frac{\partial\hat{R}(s,a,w)}{\partial w}+\sum_{(s,a)\in\tau^{i}_{2}}\frac{\partial\hat{R}(s,a,w)}{\partial w}$
$L_{2}=$ $F\Bigl{(}\sum_{(s,a)\in\tau^{i}_{1}}\hat{R}(s,a)-\sum_{(s,a)\in\tau^{i}_{2}}\hat{R}(s,a)\Bigr{)}$ (5)
$\frac{\partial L_{2}}{\partial w}=$ $F^{\prime}\Bigl{(}\sum_{(s,a)\in\tau^{i}_{1}}\hat{R}(s,a)-\sum_{(s,a)\in\tau^{i}_{2}}\hat{R}(s,a)\Bigr{)}$
$\cdot\Bigl{(}\sum_{(s,a)\in\tau^{i}_{1}}\frac{\partial\hat{R}(s,a,w)}{\partial w}-\sum_{(s,a)\in\tau^{i}_{2}}\frac{\partial\hat{R}(s,a,w)}{\partial w}\Bigr{)}$
$L_{3}=$ $\sum_{(s,a,s^{\prime})\in\tau^{i}_{1}}|\hat{Q}(s,a)-(1+\gamma\cdot\max_{a^{\prime}}\hat{Q}(s^{\prime},a^{\prime}))|+$ (6)
$\sum_{(s,a,s^{\prime})\in\tau^{i}_{2}}|\hat{Q}(s,a)-(-1+\gamma\cdot\max_{a^{\prime}}\hat{Q}(s^{\prime},a^{\prime}))|$
$\frac{\partial L_{3}}{\partial w}=$ $\sum_{(s,a,s^{\prime})\in\tau^{i}_{1}}\frac{\partial\hat{Q}(s,a)-\gamma\cdot\max_{a^{\prime}}\hat{Q}(s^{\prime},a^{\prime}))}{\partial w}-$
$\sum_{(s,a,s^{\prime})\in\tau^{i}_{2}}\frac{\partial\hat{Q}(s,a)-\gamma\cdot\max_{a^{\prime}}\hat{Q}(s^{\prime},a^{\prime}))}{\partial w}$
$L_{4}=$ $F(\sum_{(s,a,s^{\prime})\in\tau^{i}_{1}}(\hat{Q}(s,a)-\gamma\cdot\max_{a^{\prime}}\hat{Q}(s^{\prime},a)))-$ (7)
$\sum_{(s,a,s^{\prime})\in\tau^{i}_{2}}(\hat{Q}(s,a)-\gamma\cdot\max_{a^{\prime}}\hat{Q}(s^{\prime},a)))$
$\frac{\partial L_{4}}{\partial w}=$ $F^{\prime}(\sum_{(s,a,s^{\prime})\in\tau^{i}_{1}}(\hat{Q}(s,a)-\gamma\cdot\max_{a^{\prime}}\hat{Q}(s^{\prime},a)))-$
$\sum_{(s,a,s^{\prime})\in\tau^{i}_{2}}(\hat{Q}(s,a)-\gamma\cdot\max_{a^{\prime}}\hat{Q}(s^{\prime},a)))$
$\cdot$ $(\sum_{(s,a,s^{\prime})\in\tau^{i}_{2}}\frac{\partial\hat{Q}(s,a)-\gamma\cdot\max_{a^{\prime}}\hat{Q}(s^{\prime},a))}{\partial w}-$
$\sum_{(s,a,s^{\prime})\in\tau^{i}_{1}}\frac{\partial\hat{Q}(s,a)-\gamma\cdot\max_{a^{\prime}}\hat{Q}(s^{\prime},a))}{\partial w})$
Comparing the results and utilizing the fact that the loss-link function is monotonically decreasing, we get the statements in the third case from both theorems.
Appendix C Additional Experiment Results
Combing BRL with different standard Offline RL Algorithms: Here, we examine the efficiency of our method when combined with different standard offline RL algorithms training on the same dataset. We observe in Figure 5 that the learning efficiency is higher if one combines BRL with CQL when learning on the HalfCheetah medium-expert dataset. In general, the efficacy of our method is high as long as the offline RL is effective when training on the true rewards.
Figure 5: Comparison between the learning efficiency of BRL combined with different standard offline RL algorithms.
In Figure 6, we show the training logs of algorithms learning on different datasets in the main result.
In Table 2, we show the gap between the reward labels given by BRL and RM on the preferred and rejected trajectories. Note that the trajectories in the dataset are overlapped in this case.
experiment_name BRL Reward Gap RM Reward Gap
HalfCheetah-ME-(10%x2) 1.930 0.420
HalfCheetah-ME-(10%x4) 1.797 0.413
HalfCheetah-ME-(20%x2) 1.859 0.411
HalfCheetah-ME-(20%x4) 1.581 0.420
Hopper-MR-(10%x2) 1.933 0.330
Hopper-MR-(10%x4) 1.810 0.312
Hopper-MR-(20%x2) 1.876 0.317
Hopper-MR-(20%x4) 1.628 0.337
Table 2: Reward Gap by BRL and RM on preferred vs rejected trajectories